# VideoDrafter: Content-Consistent Multi-Scene Video Generation with LLM

### v1

@@ Abstract
The recent innovations and breakthroughs in diffusion models have significantly expanded the possibilities of generating high-quality videos for the given prompts. Most existing works tackle the single-scene scenario with only one video event occurring in a single background. Extending to generate multi-scene videos nevertheless is not trivial and necessitates to nicely manage the logic in between while preserving the consistent visual appearance of key content across video scenes. In this paper, we propose a novel framework, namely VideoDrafter, for content-consistent multi-scene video generation. Technically, VideoDrafter leverages Large Language Models (LLM) to convert the input prompt into comprehensive multi-scene script that benefits from the logical knowledge learnt by LLM. The script for each scene includes a prompt describing the event, the foreground/background entities, as well as camera movement. VideoDrafter identifies the common entities throughout the script and asks LLM to detail each entity. The resultant entity description is then fed into a text-to-image model to generate a reference image for each entity. Finally, VideoDrafter outputs a multi-scene video by generating each scene video via a diffusion process that takes the reference images, the descriptive prompt of the event and camera movement into account. The diffusion model incorporates the reference images as the condition and alignment to strengthen the content consistency of multi-scene videos. Extensive experiments demonstrate that VideoDrafter outperforms the SOTA video generation models in terms of visual quality, content consistency, and user preference.
Figure 1: An illustration of the input prompt and the generated multi-scene videos by using our VideoDrafter.
@@ 1 Introduction
Diffusion Probabilistic Models (DPM) have demonstrated high capability in generating high-quality images [44, 46, 12, 32, 4, 11, 40, 62, 31]. DPM approaches image generation as a multi-step sampling process, involving the use of a denoiser network to progressively transform a Gaussian noise map into an output image. Compared to 2D images, videos have an additional time dimension, which introduces more challenges when extending DPM to video domain. One typical way is to leverage pre-trained text-to-image models to produce video frames [55, 36, 16] or utilize a 3D denoiser network learnt on video data to generate a sequence of frames in an end-to-end manner [13, 43, 9, 8, 30, 2]. Despite having impressive results in the realm of text-to-video generation, most existing works focus on only single-scene videos, featuring one video event in a single background. The generation of multi-scene videos is still a problem not yet fully explored in the literature.
@@
To consolidate the idea, we present a new framework dubbed as VideoDrafter for content-consistent multi-scene video generation. Technically, VideoDrafter first transforms the input prompt into a thorough multi-scene video script by using LLM. The script for each scene consists of the descriptive prompt of the event in the scene, a list of foreground objects or persons, the background, and camera movement. VideoDrafter then identifies common entities that appear across multiple scenes and requests LLM to enrich each entity. The resultant entity description is fed into a pre-trained Stable Diffusion [40] model to produce a reference image for each entity. Finally, VideoDrafter outputs a multi-scene video via involving two diffusion models, i.e., VideoDrafter-Img and VideoDrafter-Vid. VideoDrafter-Img is dedicated to incorporating the descriptive prompt of the event and the reference images of entities in each scene as the condition to generate a scene-reference image. VideoDrafter-Vid takes the scene-reference image plus temporal dynamics of the action depicted in the descriptive prompt of the event and camera movement in the script as the inputs and produces a video clip for each scene.
The main contribution of this work is the proposal of VideoDrafter for generating content-consistent multi-scene videos. The solution also leads to the elegant views of how to use LLM to properly arrange the content of multi-scene videos and how to generate visually consistent entities across scenes, which are problems seldom investigated in the literature. Extensive experiments conducted on several public benchmarks demonstrate that VideoDrafter outperforms state-of-the-art video generation models in terms of visual quality, content consistency and user preference.
@@ 2 Related Work
Video generation is a natural extension of image generation in video domain. The early approaches, e.g., ImagenVideo [13] and Make-A-Video [43], train video diffusion models in the pixel space, resulting in high computational complexity. Following LDM in image domain, several works  [30, 2, 8] propose to exploit the sampling process in the latent feature space for video generation. These works extend the 2D UNet with transformer layers [20, 57, 58] in LDM to 3D UNet by injecting temporal self-attentions [26, 25] and/or temporal convolutions [38, 24, 27]. For instance, Video LDM [2] and AnimateDiff [8] focus on training the injected temporal layers while freezing the spatial layers to preserve the ability of the pre-trained image diffusion model. VideoFusion [30] decomposes the 3D noise into a 2D base noise shared across frames and a 3D residual noise, enhancing the correlation between frames. However, the generated videos usually have a limited time duration, typically around 16 frames. Consequently, some recent researches emerge to generate long videos by an extrapolation strategy or hierarchical architecture [9, 21, 48, 49, 60]. In addition, video editing techniques utilize the input video as a condition and generate a video by modifying the style or key object of the input video [55, 42, 36, 15, 7, 35, 6, 9, 53, 59, 50].
@@
In short, our work in this paper focuses on content-consistent multi-scene video generation. The most related work is [22], which aligns the appearance of entities across scenes through the bounding boxes provided by LLM. Ours is different in the way that we explicitly determine the appearance of entities by generating reference images, which serve as a link across scenes and effectively enhance the content consistency within a multi-scene video.
@@
Figure 2: An overview of our VideoDrafter framework for content-consistent multi-scene video generation. VideoDrafter consists of three main stages: (1) multi-scene video script generation, (2) entity reference image generation, and (3) video scene generation. In the first stage, LLM is utilized to convert the input prompt into a comprehensive multi-scene script. The script for each scene includes the descriptive prompt of the event in the scene, a list of foreground objects or persons, the background, and camera movement. We then request LLM to detail the common foreground/background entities across scenes. These entity descriptions are fed into a text-to-image (T2I) model to produce reference images in the second stage. Finally, in the third stage, VideoDrafter-Img exploits the descriptive prompt of the event and the reference images of entities in each scene as the condition to generate a scene-reference image. VideoDrafter-Vid takes the scene-reference image plus temporal dynamics of the action depicted in the descriptive prompt of the event and camera movement in the script as the inputs and produces a video clip for each scene.
3 VideoDrafter
@@
This section presents the proposed VideoDrafter framework for content-consistent multi-scene video generation. Figure 2 illustrates an overview of VideoDrafter framework, consisting of three main stages: (1) multi-scene video script generation (Sec. 3.1), (2) entity reference image generation (Sec. 3.2), and (3) video scene generation (Sec. 3.3).
@@ 3.1 Multi-Scene Video Script Generation
As depicted in Figure 2(1), VideoDrafter utilizes LLM to convert the input prompt into a comprehensive multi-scene script. In view of its high deployment flexibility and inference efficiency, we use the open-source ChatGLM3-6B model [61, 5]. The LLM is requested by a pre-defined query, “You need to envision a multi-scene video and describe each scene …”, to treat the input prompt as the theme, logically decompose the video into multiple scenes and generate a script for each scene in the following format:
@@
Next, VideoDrafter identifies the common entities, which include foreground objects or persons and background locations. To achieve this, we ask the LLM to assign the common object, person, or background the same name across scenes when generating the video script. Therefore, we strictly match the name of entities and discover the entities that appear in multiple scenes. To further improve the quality of the video script, we employ the capability of the LLM for multi-round dialogue. Specifically, we start the dialogue by asking the LLM to specify the key aspects with respect to the entity, such as “What are the aspects that should be considered when describing a photo of a young man in detail?” In the next round of dialogue, we request the LLM to describe the entity from the viewpoints of the given aspects. Moreover, the original prompt is also taken as the input to the LLM to ensure that the essential characteristics, e.g., “blue hair” of the young man, are emphasized in entity description generation.
Please note that the GPT-4 [34] can also be used for script generation, but it incurs an additional 0.12 USD for the GPT-4 API call per query. In VideoDrafter, we leverage the open-source ChatGLM3-6B and perform the inference on our devices to circumvent the need for API call. Nevertheless, the scale of ChatGLM3-6B is much smaller, resulting in unstable outcomes that may deviate from the specified script format. To alleviate this issue, we have empirically abstracted the following principles to enhance the stability of open-source LLM:
@@
    After each round of dialogue, we verify the output format. If the results are seemingly inappropriate, we re-run the entire script generation stage. Such strategy is simple to implement and does not require any additional expenses.
@@ 3.2 Entity Reference Image Generation
[…]
@@ 3.3 Video Scene Generation
VideoDrafter produces a multi-scene video by generating each scene via the diffusion models by taking the reference images, the descriptive prompt of the event and camera movement into account. This stage involves two primary components: the VideoDrafter-Img, which utilizes the descriptive prompt of the event and the reference images of entities in each scene as the condition to generate a scene-reference image, and the VideoDrafter-Vid, which employs the scene-reference image plus temporal dynamics of the action depicted in the descriptive prompt of the event and camera movement in the script as the inputs and produces a video clip for each scene.
3.3.1 VideoDrafter-Img
The VideoDrafter-Img component aims to generate a scene-reference image conditioning on the event prompt and entity reference images for each scene. To accomplish this, we remold the Stable Diffusion architecture by replacing the original attention module with a novel attention module that can handle three contexts: the text prompt, foreground reference image, and background reference image. As depicted in Figure 3, we utilize text and visual encoder of a pre-trained CLIP model to extract the sequential text feature $y_{t}\in\mathbb{R}^{L_{t}\times C_{t}}$ and local image features $y_{f}\in\mathbb{R}^{L_{f}\times C_{f}}$ and $y_{b}\in\mathbb{R}^{L_{b}\times C_{b}}$ for the prompt, foreground reference image, and background reference image, respectively. Here, $L$ and $C$ denote the length and the channels of the feature sequence. For the case of multiple foregrounds in one scene, we concatenate the features from all foreground reference images along the length dimension. Given the input feature $\bm{x}$, the outputs $\bm{z}$ of the attention are computed as
@@ 3.3.1 VideoDrafter-Img
[…]
Figure 3: A diagram illustrating the proposed attention module in the VideoDrafter-Img which takes the scene prompt and foreground/background reference images as inputs.
3.3.2 VideoDrafter-Vid
VideoDrafter-Vid is a video diffusion model that employs the scene-reference image, the action described in the prompt of the event, and camera movement in the script as the inputs. Particularly, we start by extending the Stable Diffusion model to a spatio-temporal form and replacing the original attention module with a new one that is conditioned on the scene-reference image and action category, as shown in Figure 4. Taking 400 action categories in Kinetics [3] as an action vocabulary, an indicator vector $y_{a}\in[0,1]^{400}$ is built to infer if each action in the vocabulary exists in the scene prompt and subsequently converted into feature space using a linear embedding $f$. For the scene-reference image, we use the visual encoder of CLIP to extract the image feature $y_{s}\in\mathbb{R}^{L_{s}\times C_{s}}$, which is then fed into the cross-attention operation. The original self-attention is decomposed into a spatial self-attention (Spatial SA) and a temporal self-attention (Temporal SA), which operate self-attention solely on spatial and temporal dimension, respectively, to reduce computations. Hence, given the input feature $\bm{x}$, the attention module is formulated as
@@ 3.3.2 VideoDrafter-Vid
Moreover, we further inject several temporal convolutions behind each spatial convolution into the Stable Diffusion model, to better capture temporal dependencies.
Figure 4: A diagram illustrating the proposed attention module in the VideoDrafter-Vid conditioning on the scene-reference image and the described action category.
@@ 4.1 Datasets
Our VideoDrafter framework is trained on three large-scale datasets: LAION-2B [41], WebVid-10M [1] and HD-VG-130M [52]. The LAION-5B is one of the largest text-image dataset consisting of around 5 billion text-image pairs. To train VideoDrafter-Img, We utilize a subset, namely LAION-2B, which focuses on the text prompts in English. The WebVid-10M and HD-VG-130M are the large-scale single-scene video datasets, containing approximately 10M and 130M text-video pairs, respectively. VideoDrafter-Vid is trained on the combination of WebVid-10M and a randomly chosen 20M subset from HD-VG-130M.
To evaluate video generation, we select the text prompts from three video datasets, i.e., MSR-VTT [56], ActivityNet Captions [18] and Coref-SV [22]. The first one provides the single-scene prompts, while the remaining two datasets comprise multi-scene prompts. The MSR-VTT [56] consists of 10K web video clips, each annotated with approximate 20 natural sentences. We utilize the text annotation of validation videos to serve as single-scene prompts in our evaluation. The ActivityNet Captions dataset is a multi-event video dataset designed for dense-captioning tasks. Following [22], we randomly sample 165 videos from the validation set and exploit the event captions as the multi-scene prompts. The Coref-SV is a multi-scene description dataset, which was constructed by replacing the subject of multi-scene paragraphs in Pororo-SV dataset [17, 19]. Coref-SV samples 10 episodes from the Pororo-SV dataset and replaces the subject with 10 real-world entities, resulting in 100 multi-scene prompts.

### v2

# VideoStudio: Generating Consistent-Content and Multi-Scene Videos

@@ Abstract
The recent innovations and breakthroughs in diffusion models have significantly expanded the possibilities of generating high-quality videos for the given prompts. Most existing works tackle the single-scene scenario with only one video event occurring in a single background. Extending to generate multi-scene videos nevertheless is not trivial and necessitates to nicely manage the logic in between while preserving the consistent visual appearance of key content across video scenes. In this paper, we propose a novel framework, namely VideoStudio, for consistent-content and multi-scene video generation. Technically, VideoStudio leverages Large Language Models (LLM) to convert the input prompt into comprehensive multi-scene script that benefits from the logical knowledge learnt by LLM. The script for each scene includes a prompt describing the event, the foreground/background entities, as well as camera movement. VideoStudio identifies the common entities throughout the script and asks LLM to detail each entity. The resultant entity description is then fed into a text-to-image model to generate a reference image for each entity. Finally, VideoStudio outputs a multi-scene video by generating each scene video via a diffusion process that takes the reference images, the descriptive prompt of the event and camera movement into account. The diffusion model incorporates the reference images as the condition and alignment to strengthen the content consistency of multi-scene videos. Extensive experiments demonstrate that VideoStudio outperforms the SOTA video generation models in terms of visual quality, content consistency, and user preference. Source code is available at https://github.com/FuchenUSTC/VideoStudio.
Figure 1: An illustration of prompt and multi-scene video generation by VideoStudio.
@@ 1 Introduction
Diffusion Probabilistic Models (DPM) have demonstrated high capability in generating high-quality images [48, 50, 15, 36, 7, 16, 44, 68, 35]. DPM approaches image generation as a multi-step sampling process, involving the use of a denoiser network to progressively transform a Gaussian noise map into an output image. Compared to 2D images, videos have an additional time dimension, which introduces more challenges when extending DPM to video domain. One typical way is to leverage pre-trained text-to-image models to produce video frames [59, 40, 20] or utilize a 3D denoiser network learnt on video data to generate a sequence of frames in an end-to-end manner [14, 47, 12, 11, 34, 3]. Despite having impressive results in the realm of text-to-video generation, most existing works focus on only single-scene videos, featuring one event in a single background. The generation of multi-scene video is still a problem not yet fully explored in the literature.
@@
To consolidate the idea, we present a new framework dubbed as VideoStudio for consistent-content and multi-scene video generation. Technically, VideoStudio first transforms the input prompt into a thorough multi-scene video script by using LLM. The script for each scene consists of the descriptive prompt of the event in the scene, a list of foreground objects or persons, the background, and camera movement. VideoStudio then identifies common entities that appear across multiple scenes and requests LLM to enrich each entity. The resultant entity description is fed into a pre-trained Stable Diffusion [44] model to produce a reference image for each entity. Finally, VideoStudio outputs a multi-scene video via involving two diffusion models, i.e., VideoStudio-Img and VideoStudio-Vid. VideoStudio-Img is dedicated to incorporating the descriptive prompt of the event and the reference images of entities in each scene as the condition to generate a scene-reference image. VideoStudio-Vid takes the scene-reference image plus temporal dynamics of the action depicted in the descriptive prompt of the event and camera movement in the script as the inputs and produces a video clip for each scene.
The main contribution of this work is the proposal of VideoStudio for generating consistent-content and multi-scene videos. The solution also leads to the elegant views of how to use LLM to properly arrange content of multi-scene videos and how to generate visually consistent entities across scenes, which are problems seldom investigated in literature. Extensive experiments conducted on public benchmarks demonstrate that VideoStudio outperforms SOTA video generation models in terms of visual quality, content consistency and user preference.
@@ 2 Related Work
Video generation is a natural extension of image generation in video domain. The early approaches, e.g., ImagenVideo [14] and Make-A-Video [47], train video diffusion models in the pixel space, resulting in high computational complexity. Following LDM in image domain, several works  [34, 3, 11, 70, 6] propose to exploit the sampling process in the latent feature space for video generation. These works extend the 2D UNet with the transformer layers [23, 62, 63] to 3D UNet by injecting temporal self-attentions [28, 29] and/or temporal convolutions [30, 31]. For instance, Video LDM [3] and AnimateDiff [11] focus on training the injected temporal layers while freezing the spatial layers to preserve the ability of the pre-trained image diffusion model. VideoFusion [34] decomposes the 3D noise into a 2D base noise shared across frames and a 3D residual noise, enhancing the correlation between frames. However, the generated videos usually have a limited time duration, typically around 16 frames. Consequently, some recent researches emerge to generate long videos by an extrapolation strategy or hierarchical architecture [12, 25, 52, 53, 66]. In addition, video editing techniques utilize the input video as a condition and generate a video by modifying the style or key object of the input video [59, 46, 40, 18, 10, 39, 9, 12, 57, 65, 54].
@@
In short, our work in this paper focuses on consistent-content and multi-scene video generation. The most related work is [26], which aligns the appearance of entities across scenes through the bounding boxes provided by LLM. Ours is different in the way that we explicitly determine the appearance of entities by generating reference images, which serve as a link across scenes and effectively enhance the content consistency within a multi-scene video.
@@
Figure 2: An overview of our VideoStudio framework for consistent-content and multi-scene video generation. VideoStudio consists of three main stages: (1) multi-scene video script generation, (2) entity reference image generation, and (3) video scene generation. In the first stage, LLM is utilized to convert the input prompt into a comprehensive multi-scene script. The script for each scene includes the descriptive prompt of the event in the scene, a list of foreground objects or persons, the background, and camera movement. We then request LLM to detail the common foreground/background entities across scenes. These entity descriptions are fed into a text-to-image (T2I) model to produce reference images in the second stage. Finally, in the third stage, VideoStudio-Img exploits the descriptive prompt of the event and the reference images of entities in each scene as the condition to generate a scene-reference image. VideoStudio-Vid takes the scene-reference image plus temporal dynamics of the action depicted in the descriptive prompt of the event and camera movement in the script as the inputs and produces a video clip for each scene.
3 VideoStudio
@@
This section presents the proposed VideoStudio framework for consistent-content and multi-scene video generation. Figure 2 illustrates an overview of VideoStudio framework, consisting of three main stages: (1) multi-scene video script generation (Sec. 3.1), (2) entity reference image generation (Sec. 3.2), and (3) video scene generation (Sec. 3.3).
@@ 3.1 Multi-Scene Video Script Generation
As depicted in Figure 2(1), VideoStudio utilizes LLM to convert the input prompt into a comprehensive multi-scene script. In view of its high deployment flexibility and inference efficiency, we use the open-source ChatGLM3-6B model [67, 8]. The LLM is requested by a pre-defined query, “You need to envision a multi-scene video and describe each scene …”, to treat the input prompt as the theme, logically decompose the video into multiple scenes and generate a script for each scene in the following format:
@@
Next, VideoStudio identifies the common entities, which include foreground objects or persons and background locations. To achieve this, we ask the LLM to assign the common object, person, or background the same name across scenes when generating the video script. Therefore, we strictly match the name of entities and discover the entities that appear in multiple scenes. To further improve the quality of the video script, we employ the capability of the LLM for multi-round dialogue. Specifically, we start the dialogue by asking the LLM to specify the key aspects with respect to the entity, such as “What are the aspects that should be considered when describing a photo of a young man in detail?” In the next round of dialogue, we request the LLM to describe the entity from the viewpoints of the given aspects. Moreover, the original prompt is also taken as the input to the LLM to ensure that the essential characteristics, e.g., “blue hair” of the young man, are emphasized in entity description generation.
Please note that the GPT-4 [38] can also be used for script generation, but it incurs an additional 0.12 USD for the GPT-4 API call per query. In VideoStudio, we leverage the open-source ChatGLM3-6B and perform the inference on our devices to circumvent the need for API call. Nevertheless, the scale of ChatGLM3-6B is much smaller, resulting in unstable outcomes that may deviate from the specified script format. To alleviate this issue, we have empirically abstracted the following principles to enhance the stability of open-source LLM:
@@
    After each round of dialogue, we verify the output format. If the results are seemingly inappropriate, we re-run the entire script generation stage. Such strategy is simple to implement without requiring any additional expenses.
@@ 3.2 Entity Reference Image Generation
In the second stage of VideoStudio, we unify the visual appearance of common entities by explicitly generating a reference image for each entity. The reference images act as the link to cohere the content across scenes. We achieve this by first feeding the entity description into a pre-trained Stable Diffusion model for text-to-image generation. Then, we employ the U2-Net [41] model for salient object detection, and segment the foreground and background areas in each resultant image. By utilizing the segmentation masks, we can further remove the background pixels from the foreground reference image and vice versa, in order to prevent the interference between the foreground and background visual contents in the reference images.
@@ 3.3 Video Scene Generation
VideoStudio produces a multi-scene video by generating each scene via the diffusion models by taking the reference images, the descriptive prompt of the event and camera movement into account. This stage involves two primary components: the VideoStudio-Img, which utilizes the descriptive prompt of the event and the reference images of entities in each scene as the condition to generate a scene-reference image, and the VideoStudio-Vid, which employs the scene-reference image plus temporal dynamics of the action depicted in the descriptive prompt of the event and camera movement in the script as the inputs and produces a video clip for each scene.
3.3.1
The VideoStudio-Img component aims to generate a scene-reference image conditioning on the event prompt and entity reference images for each scene. To accomplish this, we remold the Stable Diffusion architecture by replacing the original attention module with a novel attention module that can handle three contexts: the text prompt, foreground reference image, and background reference image. As depicted in Figure 3, we utilize text and visual encoder of a pre-trained CLIP model to extract the sequential text feature $y_{t}\in\mathbb{R}^{L_{t}\times C_{t}}$ and local image features $y_{f}\in\mathbb{R}^{L_{f}\times C_{f}}$ and $y_{b}\in\mathbb{R}^{L_{b}\times C_{b}}$ for the prompt, foreground reference image, and background reference image, respectively. Here, $L$ and $C$ denote the length and the channels of the feature sequence. For the case of multiple foregrounds in one scene, we concatenate the features from all foreground reference images along the length dimension. Given the input feature $\bm{x}$, the outputs $\bm{z}$ of the attention are computed as
@@
where CA1 and SA are the cross-attention and self-attention modules, respectively, in the original Stable Diffusion architecture. We add two additional cross-attention modules, CA2 and CA3, which leverage the guidance provided by entity reference images. Moreover, we propose to optimize the parameters of CA2 and CA3 while freezing the other parts of the network.
Figure 3: Diagram illustrations of (a) attention module in the VideoStudio-Img which takes the scene prompt and foreground/background reference images as the inputs and (b) attention module in the VideoStudio-Vid conditioning on the scene-reference image and the described action category.
3.3.2
The VideoStudio-Vid is a video diffusion model that employs the scene-reference image, the action described in the prompt of the event, and camera movement in the script as the inputs. Particularly, we start by extending the Stable Diffusion model to a spatio-temporal form and replacing the original attention module with a new one that is conditioned on the scene-reference image and action category, as shown in Figure 3. Taking 400 action categories in Kinetics [4] as an action vocabulary, an indicator vector $y_{a}\in[0,1]^{400}$ is built to infer if each action in the vocabulary exists in the scene prompt and subsequently converted into feature space using a linear embedding $f$. For the scene-reference image, we use the visual encoder of CLIP to extract the image feature $y_{s}\in\mathbb{R}^{L_{s}\times C_{s}}$, which is then fed into the cross-attention operation. The original self-attention is decomposed into a spatial self-attention (Spatial SA) and a temporal self-attention (Temporal SA), which operate self-attention solely on spatial and temporal dimension, respectively, to reduce computations. Hence, given the input feature $\bm{x}$, the attention module is formulated as
@@
Moreover, we further inject several temporal convolutions behind each spatial convolution into the Stable Diffusion architecture, to better capture temporal dependencies in image-to-video generation.
@@ 4.1 Datasets
Our VideoStudio framework is trained on three large-scale datasets: LAION-2B [45], WebVid-10M [1] and HD-VG-130M [56]. The LAION-5B is one of the largest text-image dataset consisting of around 5 billion text-image pairs. To train VideoStudio-Img, We utilize a subset, namely LAION-2B, which focuses on the text prompts in English. The WebVid-10M and HD-VG-130M are the large-scale single-scene video datasets, containing approximately 10M and 130M text-video pairs, respectively. VideoStudio-Vid is trained on the combination of WebVid-10M and a randomly chosen 20M subset from HD-VG-130M.
To evaluate video generation, we select the text prompts from three video datasets, i.e., MSR-VTT [61], ActivityNet Captions [22] and Coref-SV [26]. The first one provides the single-scene prompts, while the remaining two datasets comprise multi-scene prompts. The MSR-VTT consists of 10K web video clips, each annotated with approximate 20 natural sentences. We utilize the text annotation of validation videos to serve as single-scene prompts in our evaluation. The ActivityNet Captions dataset is a multi-event video dataset designed for dense-captioning tasks. Following [26], we randomly sample 165 videos from the validation set and exploit the event captions as the multi-scene prompts. The Coref-SV is a multi-scene description dataset, which was constructed by replacing the subject of multi-scene paragraphs in Pororo-SV dataset [21, 24]. Coref-SV samples 10 episodes from the Pororo-SV dataset and replaces the subject with 10 real-world entities, resulting in 100 multi-scene prompts.